1) Let f(x) = π − √7x − 9. Use the definition of the derivative to find f′(x).
2) Differentiate the following function
g(x)=e^-5x + (1/x −tanx)^3 (√x^2 +1)
1
Expert's answer
2020-12-15T07:06:24-0500

1) From the definition of derivative of a function "f(x)" we know that,

"f'(x)=\lim_{h\to0} [f(x+h)-f(x)]/h" ....(1)

According to the problem given that,

"f(x)=\pi-\sqrt{(7x-9)}"

So,"f(x+h)=\pi-\sqrt{7(x+h)-9}"

Therefore from (1),we have

"f'(x)=\lim_{h\to0}[f(x+h)-f(x)]/h" "=lim_{h\to0}[[\pi-\sqrt{7(x+h)-9}]-[\pi-\sqrt{7x-9}]]/h"

"=lim_{h\to0}[\sqrt{7x-9}-\sqrt{7(x+h)-9}]/h"

"=lim_{h\to0}[(\sqrt{7x-9}-\sqrt{7(x+h)-9})×(\sqrt{7x-9}+\sqrt{7(x+h)-9})]\div[h×(\sqrt{7x-9}+\sqrt{7(x+h)-9})]"

"=lim_{h\to0}[(7x-9)-(7(x+h)-9)]\div[h×(\sqrt{7x-9}+\sqrt{7(x+h)-9})]"

"=lim_{h\to0}(-7h)/[h×(\sqrt{7x-9}+\sqrt{7(x+h)-9}]"

"=lim_{h\to0}(-7)/[\sqrt{7x-9}+\sqrt{7(x+h)-9}]"

"=(-7)/[\sqrt{7x-9}+\sqrt{7(x+0)-9}]"

"=(-7)/[2×\sqrt{7x-9}]"

Therefore required "f'(x)=(-7)/[2×\sqrt{7x-9}]"




2)Given that

"g(x)=e^{-5x}+[(1/x)-tanx]^3.\sqrt{x^2+1}"

"g'(x)=d/dx(e^{-5x})+d/dx[[(1/x)-tanx]^3.\sqrt{x^2+1}]"

"=-5.e^{-5x}+[(1/x)-tanx]^3.d/dx[\sqrt{x^2+1}]+\sqrt{x^2+1}.d/dx[(1/x)-tanx]^3"

"=-5.e^{-5x}+[(1/x)-tanx]^3.[1/2{\sqrt{x^2+1}}].d/dx(x^2+1)+\sqrt{x^2+1}×3[(1/x)-tanx]^2.d/dx[(1/x)-tanx]"

"=-5.e^{-5x}+[(1/x)-tanx]^3.[x/\sqrt{x^2+1}]+3\sqrt{x^2+1}×[(1/x)-tanx]^2.[(-1/x^2)-sec^2x]"